# RAS: Retrieval-And-Structuring for Knowledge-Intensive LLM Generation

### v1

@@ Appendix F Qualitative Examples
Appendix G Hyperparameter Study
@@
| Theme Classifier | |
| Encoder learning rate | {1e-6, 1e-5, 2e-5, 1e-4, 2e-4} |
| Classifier learning rate | {1e-6, 1e-5, 2e-5, 1e-4, 2e-4, 5e-4, 1e-3} |
| Batch size | {4, 8, 16, 32, 64} |
| Hidden layers | {1, 2, 3} |
| Hidden dim | {256, 512, 1024} |
| Dropout rate | {0.1, 0.2, 0.3} |
| Distribution Shifter | |
| Batch size | {4, 8, 16, 32, 64} |
| Learning rate | {0.0001, 0.0005, 0.001, 0.005} |
| Hidden layers | {[256], [512], [512, 256], [1024, 512, 256]} |
| Dropout rate | {0.1, 0.15, 0.2, 0.25, 0.3} |
| Optimizer | {Adam, SGD, RMSprop, AdamW} |
| Loss function | {KLDiv, JSD, MSE, Wasserstein, CrossEntropy} |
@@
Appendix H Notation Table
| Notation | Description |
| General | |
| Main input query | |
| Generated answer | |
| Complete document corpus | |
| Theme-indexed document corpus | |
| Action Planning | |
| Action planner model | |
| Plan generated at iteration | |
| Subquery generated at iteration | |
| Graph information (triple list) extracted at iteration | |
| Theme-Scoped Retrieval | |
| Theme classifier function | |
| Theme distribution shifter function | |
| Dense retriever function | |
| Posteriori probability distribution of query | |
| Expected document theme distribution | |
| Theme-specific sub-corpus at iteration | |
| Retrieved text passages at iteration | |
| Number of documents selected in theme scoping | |
| Number of passages selected in dense retrieval | |
| Knowledge Structuring | |
| Text-to-triple conversion model | |
| Text encoder for node/edge attributes | |
| Graph neural network encoder | |
| Knowledge graph at iteration | |
| Final query-specific knowledge graph | |
| Node and edge sets of graph at iteration | |
| Node and edge sets of final graph | |
| Attribute vector of node | |
| Attribute vector of edge | |
| Answer Generation | |
| Answerer model | |
| Model-generated answer before retrieval | |

### v2

@@ Appendix G Qualitative Examples
Appendix H Hyperparameter Study
@@
Appendix I Broader Impacts, Safeguards, and Assets
Broader Impacts. The RAS framework presents several promising impacts for both research and practical applications. By transforming unstructured retrieved content into structured, query-specific knowledge graphs, RAS improves reasoning transparency and factual reliability—qualities that are especially valuable in education, scientific research, and technical writing. The ability to explicitly organize retrieved knowledge into interpretable structures also opens up new possibilities for human-in-the-loop AI systems, enabling users to trace and verify model reasoning paths.
However, several ethical considerations remain. RAS’s reliance on structured knowledge representations means its effectiveness may be limited in low-resource languages or domains with sparse data. Additionally, biases present in training data can be preserved or even amplified in structured forms. Addressing these concerns will require rigorous evaluation across diverse user groups and careful curation of source corpora. Promoting fairness, transparency, and inclusivity should be core design goals for future RAS-based systems.
Safeguards. To mitigate potential harms, we propose several safeguards during the development and deployment of RAS-based systems. First, developers should routinely audit training and retrieval corpora for biases and factual inconsistencies, and apply data augmentation or counterfactual generation techniques where appropriate. Second, regular evaluations should be conducted across a diverse range of tasks, languages, and user demographics to assess any unintended performance disparities. Third, when applied in sensitive domains, human-in-the-loop oversight should be incorporated to validate the structured knowledge outputs and final generations. Lastly, transparency measures, such as providing users with access to intermediate structured knowledge (e.g., the constructed knowledge graph), can help foster trust and enable error analysis.
Assets Used. To support the development and evaluation of the RAS framework, we utilized a wide range of publicly available datasets, open-source tools, and pretrained models. All datasets were selected based on relevance to knowledge-intensive tasks, including question answering and long-form generation, and were licensed for academic research or distributed under permissive open-source terms such as MIT, Apache 2.0, and CC BY. Our training and evaluation relied on benchmark datasets such as HotpotQA, ASQA, TriviaQA, and ARC, while structured knowledge was derived using models trained on WikiOfGraph. We used open-source infrastructure including FAISS for retrieval, vLLM for inference optimization, and Graph LLM as our model backbone. For baseline comparisons, we incorporated implementations from Self-RAG, RPG, and ALCE. Closed-source models like Claude-3.5-Sonnet were used for performance comparison and triple extraction under API constraints. Full licensing and usage details are provided in Table 8.
| Asset | Source / URL | License | Usage |
| HotpotQA | https://hotpotqa.github.io/ | MIT | Base QA dataset |
| HotpotQA-SUBQ | We constructed from HotpotQA | MIT (inherited) | Subquery training dataset |
| ASQA | https://github.com/google-research/language/tree/master/language/asqa | Apache 2.0 | Training and evaluation |
| Arc-Easy | https://allenai.org/data/arc | AI2 Research License | Training (subset) |
| TriviaQA | https://nlp.cs.washington.edu/triviaqa/ | CC BY-SA 4.0 | Evaluation |
| PopQA | https://huggingface.co/datasets/akariasai/PopQA | CC BY 4.0 | Evaluation |
| 2WikiMultihopQA | https://github.com/Alab-NII/2wikimultihop | Apache 2.0 | Evaluation |
| PubHealth | https://huggingface.co/datasets/bigbio/pubhealth | MIT | Evaluation |
| ARC-Challenge | https://allenai.org/data/arc | AI2 Research License | Evaluation |
| ELI5 | https://huggingface.co/datasets/eli5 | Apache 2.0 | Evaluation |
| Wikipedia 2018/2020 | https://github.com/facebookresearch/atlas | Creative Commons Public Licenses | Knowledge source for retrieval |
| WikiOfGraph | https://github.com/daehuikim/WikiOFGraph | CC BY 4.0 | Triple model training |
| Claude-3.5-Sonnet | https://www.anthropic.com | Proprietary | Default closed source model |
| LLaMA-2/3 | https://ai.meta.com/llama | Meta LLaMA Community License | Default open source model |
| LLaMA-3.2-3B | Derived from Meta LLaMA | Meta LLaMA Community License | Triple extraction |
| Flan-T5-Large | https://huggingface.co/google/flan-t5-large | Apache 2.0 | Triple extractor |
| Graph LLM | Modified from https://github.com/XiaoxinHe/G-Retriever | Research use | Model backbone |
| Faiss | https://github.com/facebookresearch/faiss | MIT | Dense retrieval |
| vLLM | https://github.com/vllm-project/vllm | Apache 2.0 | Optimized inference |
| Sentence-BERT | https://www.sbert.net | Apache 2.0 | Graph node/edge embedding |
| Self-RAG Code | https://github.com/AkariAsai/self-rag | MIT | Baseline eval |
| RPG Code | https://github.com/haruhi-sudo/RPG | Research use | Baseline eval |
| ALCE Code | https://github.com/princeton-nlp/ALCE | MIT | Eval metrics |